Sol Summary – April 15th

Crew 245 Sol Summary Report 15-04-2021

Sol: 5

Summary Title: What is Nominal? / Emergency EVA Return

Author’s name: Shayna Hume

Mission Status: Executive Officer

Sol Activity Summary:

· 07:30-08:30: Blueberry pancakes with Nutella!

· 08:30-09:30: Prep for EVA & filming of prep.

· 09:30-11:15: Crew #2 (Hume, Hernandez, Hariharan) left for EVA #4. On this EVA we intended to visit site #2 – at GPS coordinates on the edge of an interesting geologic ridge. We not only managed to identify a possible emergency weather shelter, but sampled three different locations for soil analysis. We noticed clouds getting darker and started heading back. By the time we got to the vehicles, there was significant (~35mph) wind and we felt hurried by inclement emergency conditions. We got back about an hour ahead of schedule but safely, and felt proud of our handling of the situation.

· 12:00-14:00: Spaghetti & Soup for lunch. Hot chocolate on the side!

· 14:00-17:00: After taking an extra long lunch we separated for work. Commander and HSO went to Science Dome for work, Crew Scientist and XO began to work on soil sampling, Engineer and Botanist were working on misc in the GreenHab and on reports.

· 17:00-19:00: Fried rice for dinner! Attempting to send reports at 7p on the dot today to spend some quality time as a crew offline.

Look Ahead Plan: Tomorrow is EVA #5, where Crew #1 will go to a geographic gully location (site #3). At this time, we are beginning to maintain the botany experiments, move onto thicker gloves for the dexterity study, and begin to analyze the soil from the EVA’s.

Anomalies in work: Emergency weather conditions forced an early EVA return!

Weather: Warm in daytime, around 40 F at nighttime. Extremely high winds 11:00 am onwards.

Crew Physical Status: No issues.

EVA: EVA #4, with EVA crew #2. Report incoming.

Reports to be filed: Operations Report, EVA Report, EVA Plan, Science Report, Commander’s Report

Support Requested: n/a

Sol Summary Report Apr 14th

Crew 245 Sol Summary Report 14APR2021

Sol: 4

Summary Title: Exploring the Unknown / First Real EVA

Author’s name: Shayna Hume

Mission Status: Executive Officer

Sol Activity Summary:

· 07:30-08:30: Rehydrated strawberries over oatmeal by Sous Chef Shayna!

· 08:30-09:00: Prep for EVA.

· 09:00: Crew #1 (Ettlin, Dickstein, Coultrup) left for EVA #3 – the first real science EVA. There were significant radio problems since we did not know we would lose comms with distances we had mapped out, however we were able to come up with multiple options in the trade space of 1) new, closer location, and 2) confirmation of being able to go out of radio-distance. The EVA then went very successfully! Excellent handling of the problem.

· 09:00-12:00: Crew Botanist set up Martian regolith botany, Crew Engineer worked on several misc. problems, and XO wrote reports… as normal.

· 12:00: Working on veggie empanadas for lunch!

· 12:00-12:30: Crew Engineer emptied out the sewage tank.

· 12:30-17:00: Misc work throughout afternoon.

· 17:00-19:00: Dinner cooking and report writing. Mac and Cheese for Dinner!

Look Ahead Plan: Tomorrow is our second science EVA, for Crew #2 (Hernandez, Hariharan, Hume). Looking to get all botany experiments underway by tomorrow morning. May have more time for breathing finally and other crew activities.

Anomalies in work: Lost comms with EVA crew when they went at a distance. Turned out to be a non-issue.

Weather: Warm in daytime, around 40 F at nighttime. Clear weather. Heavy wind after noontime.

Crew Physical Status: No issues. Dry noses all around.

EVA: EVA #3, first science EVA with EVA crew #1. Report incoming.

Reports to be filed: Operations Report, EVA Report, EVA Plan, Science Report, Commander’s Report

Support Requested: n/a

Sol Summary – April 13th

Crew 245 Sol Summary Report 13-04-2021

Sol: 3

Summary Title: First day in Sim

Author’s name: Shayna Hume

Mission Status: Executive Officer

Sol Activity Summary:

· Woke up to “Good Morning” by Kanye…. again

· Early morning hours: Strawberry, blueberry, and chocolate chip pancakes! Chef Julio is unreal.

· Late morning hours: Hurried for a 9am first EVA. From 9-12, we did our two practice EVA’s in two crews. Crew 1: Hume, Coultrup, Ettlin. Crew 2: Dickstein, Hariharan, Hernandez. First words outside: “We go to Mars together.” Second EVA ended early due to comms breakdown.

· Noon: separated for a variety of activities. Crew Scientist doing botany. HSO starting a botany kit as well. Report-writing for the rest.

· Lunch: Chef Julio preparing quesadillas! Corn tortillas from scratch! It was absolutely wonderful.

· Afternoon: Scientist & Botanist in Greenhab, HSO and Commander working on Documentary, Engineer and XO on misc.

· Late Afternoon: Scientist & Botanist in Greenhab, Engineer grabbing people for dexterity task, HSO & XO answering misc. calls for help, Commander videoing intro videos for our mockumentary.

· Dinner: Beef tacos!

Look Ahead Plan: Our first science mission is tomorrow morning. In addition, we will be completing the second day of dexterity testing in spacesuit gloves. Misc. botany will be happening, as well as the first soil sample analysis for geology and astrobiology. We are now living fully in sim and beginning to work towards making our workload each day sustainable.

Anomalies in work: Comms went dead on EVA, ended early. Dead charge only.

Weather: Warm in daytime, around 40 F at nighttime. Clear weather. Expected rain did not fall.

Crew Physical Status: No major issues. HSO Report coming in for small aid kit use.

EVA: Two practice EVA’s. First one had several equipment issues, completed as nominal however. Second had comms issue. EVA Reports incoming.

Reports to be filed: HSO Report, Operations Report, EVA Report, EVA Plan, Science Report

Support Requested: We would like to abstain from sending 8 daily photos, as it’s taking up our whole internet. We uploaded photos over lunch and it virtually drained the entire internet for the day. A single photo may be more feasible. Please confirm a single photo of the day will do for now. We are running out just by sending this.

Shayna Hume, Executive Officer
Red Planet People – MDRS Crew 245 “Team Patamars”
To Mars and Beyond – For All!

Sol Summary – April 12th

Crew 245 Sol Summary Report 12-04-2021

Sol: 2

Summary Title: Before the Simulation

Author’s name: Shayna Hume

Mission Status: Executive Officer

Sol Activity Summary:

– Woke up to “Good Morning” by Kanye
– Early morning hours: exploring the kitchen (no can opener at the time), making breakfasts and coffees, and talking about our thoughts on the start of the sim
– Late morning hours: met up with Dr. Rupert, began to do our trainings. Had an excellent and lively discussion. Felt confident in our pre-knowledge for the station.
– Noon: took the training quiz and passed as a team.
– Late lunch: Chef Julio prepared spiced potato au gratin and potato soup. We mixed the two together and it was wonderful. Running a bit late to meet with Dr. Rupert. Took photos of a group meal for GSI!
– Mid-afternoon: Tarped the tunnels between the hab buildings quickly as a team. Continued to our atv and rover training – what a blast! Do not use the kill switch. Aside: my (Hume) callsign is now “Killswitch.” I couldn’t be more pleased. After this, split while Dr. Rupert went to help/look for Atila. Tried on flight and spacesuits, figured out helmets. Got some great photos!
– Late afternoon: As we wait for formal last trainings with Dr. Rupert, each of us took on the below tasks:
Crew Engineer and Commander: Heater (filter replacement pending) & operations report
HSO and Commander: HSO Inspection
XO: Report-writing
Scientist and Botanist: Cleaning the lab.

– Evening: Meeting with Shannon to review spacesuits, assignments, and radio protocols while in EVA. Took photos, began dinner, sent this report.

Look Ahead Plan: We intend to begin sim overnight. We will be ready to begin sim just before we retire for the evening so this makes practical sense. Today was an exciting day, because although we aren’t in sim, we are 90% of the way there. The main difference is that we are still going outside at this time. We are eating in sim, using internet in sim, and working together in sim. We are ready to begin EVA’s tomorrow. It’s beyond exciting to be about to start our first analog mission, and the energy is definitely running through the crew. We have a vision of being able to work together as a crew many times in future, and this is the precipice of beginning that journey.

Anomalies in work: n/a

Weather: Warm in daytime, around 40 F at nighttime. Clear weather.

Crew Physical Status: No issues. HSO First Day Report coming in.

EVA: n/a

Reports to be filed: HSO First Report, Operations, EVA Plan

Support Requested: n/a

Sol Summary – April 11th

Crew 245 Sol Summary Report 11-04-2021

Sol: 1

Summary Title: Arrival at the Station

Author’s name: Shayna Hume

Mission Status: Executive Officer

Sol Activity Summary: Arrive at Station. Because of the late hour, the priority was to unpack and learn the essentials of living in the station safely. We settled into our staterooms. The Crew Engineer spent time with the station Director learning how the toilet operates and flushing old water out of the system. Commander and XO spent time preparing initial emails and reports.

The primary challenge that arose was an uncertain smell near the Crew Scientist’s stateroom. It was found that it was sewage gas and the flushing of the toilet took care of the issue. A full operations report on Sol 2 is following. Note: we are not yet in simulation.

Look Ahead Plan: On Sol 2, we will be beginning the training, taking the quiz, and assisting with last-minute repairs needed for the Hab, including putting up the tarps for between buildings. Before the end of the day, we intend to

Anomalies in work: n/a

Weather: Warm in the daytime, around 40 F at nighttime. Clear weather.

Crew Physical Status: No issues. HSO Report for before the sim coming on Sol 2.

EVA: n/a

Reports to be filed: n/a

Support Requested: n/a

Shayna Hume, Executive Officer
Red Planet People – MDRS Crew 245 "Team Patamars"
To Mars and Beyond – For All!

Sol Summary – March 13th

Crew 223 Sol Summary Report 13-03-2020

Sol: 12 Summary Title: That’s all folk’s

Author’s name: Aurélien Mure – Crew Commander

Mission Status: crew out of sim

Sol Activity Summary:

The crew woke up at 7:50am. The sport session began at 8:00am. Members are struggling to sleep, that is why we moved the sport session, knowing that the weather did not allow us to go outside in the morning. We did all the cleaning work on the hab, the GreenHab, the science dome and the RAM. As we had to Unmount our LOAC and MegaAres, the crew worked on the Human Factor experiment in the late morning.

We went out with a lot of wind and after 4 weeks of hard labour, the weather station fell down one hour before we were supposed to unmount it. but no worries, nothing is broken. The EVA was not so easy because of the wind.

We washed all the materials and we went out of sim. After a long debriefing on how we have lived the mission, we were ready to cook for tonight !

Look Ahead Plan : Unmount the outdoor experiments.

Anomalies in work: None.

Weather: rainy in the morning, 4°C at 8:30am then cloudy and windy in the afternoon.

Crew Physical Status: All crew members safe and sound.

EVA: LOAC and MegaAres unmounted

Reports to be filed: Journalist report, Operations report, Green Hab report, EVA report, Astronomy report, Scientist, food inventory, Mission summary

Support Requested: None

Sol Summary – March 12th

Crew 223 Sol Summary Report 12-03-2020

Sol: 11 Summary Title: Strange feelings

Author’s name: Aurélien Mure – Crew Commander

Mission Status: Last full day of Sim

Sol Activity Summary:

The crew woke up at 6:50am. The sport session began at 7:00am. The end is close, a good team spirit was felt during the session. For the EVA, the crew explored the extreme North region of the campus. They were satisfied with what they have found. Red soil was everywhere and it made Marion happy. During the EVA, the three members in the hab prepared a joke for Florian. Indeed, they put a Jalapeño pepper inside his bread for lunch. This was a really funny moment, and with the bad news from the outside world it is good to share this. Then we talked a lot about the Covid19 consequences on the end of our year. Our school has been officially closed and those who were supposed to go to Costa Rica, Guyane or Italy are thinking about alternatives.

We stay really close from each other, which helps us to keep our spirit up!

Look Ahead Plan : Un-mount the outdoor experiments.

Anomalies in work: None.

Weather: cloudy in the morning, 0°C at 7:30am then cloudy and sunny in the late afternoon, 14°C at 7:00pm

Crew Physical Status: All crew members safe and sound.

EVA: LOAC and MegaAres batteries changed, Northern region explored

Reports to be filed: Journalist report, Operations report, EVA request, Green Hab report, EVA report, Astronomy report.

Support Requested: None

Sol Summary – March 11th

Crew 223 Sol Summary Report 11-03-2020

Sol: 10 Summary Title: Mars is safer than Earth

Author’s name: Aurélien Mure – Crew Commander

Mission Status: Routine

Sol Activity Summary:

The crew woke up at 6:50am. The sport session began at 7:00am. A light rain and the night outside were telling us warmly to go back to bed, but quickly we thought about how lucky we are to be here. Without EVA planned for the morning because of the rain, the crew started to prepare the end of the mission, thinking about packing our experiments and writing reports about the science we did during the mission. For the afternoon, Blandine and Florian went out 30min to change the batteries and data from our experiments.

We have had news about the Covid19 at 4:00 pm.

It is strange to see it from here, we are really like martians on this.

Look Ahead Plan : Unmount the outdoor experiments.

Anomalies in work: None.

Weather: rainy in the morning, then cloudy and sunny in the late afternoon, 12°C at 5:00pm

Crew Physical Status: All crew members safe and sound.

EVA: LOAC and MegaAres batteries changed.

Reports to be filed: Journalist report, Operations report, EVA request, Green Hab report, EVA report, Astronomy report.

Support Requested: None

Sol Summary – March 10th

Crew 223 Sol Summary Report 10-03-2020

Sol: 9 Summary Title: Tired crew

Author’s name: Aurélien Mure – Crew Commander

Mission Status: Routine

Sol Activity Summary:

The crew woke up at 6:50am. The sport session began at 7:00am. A few members of the crew have had a difficult night, and we have felt it during the day. Some of us took some rest after the EVA or at the beginning of the afternoon.

The EVA was a long range walk into candor, to go as far as possible. We have discovered new landscapes and really Martian Candor Chasma like.

We went through our classical science afternoon finishing with our Human Factor experiment and our relaxation and cohesive activity time.

Look Ahead Plan : Starting to think about packing our experiments.

Anomalies in work: None.

Weather: cloudy, 0°C at 7:00am, 12°C at 7:00pm

Crew Physical Status: All crew members safe and sound.

EVA: LOAC and MegaAres batteries changed. Candle Chasma exploration exploration

Reports to be filed: Journalist report, Operations report, EVA request, Green Hab report, EVA report, Astronomy report, Science report,

Support Requested: None

Sol Summary – March 9th

Crew 223 Sol Summary Report 09-03-2020

Sol: 8 Summary Title: Ambitious EVA

Author’s name: Aurélien Mure – Crew Commander

Mission Status: Science is running

Sol Activity Summary:

Let’s go back to normal. Alarm at 6:45 and sport session at 7:00. After a rest day, it is difficult to warm up! After the common breakfast, the EVA team prepared itself. They chose to go to Copernicus with the ATV’s. First problem, two of them refused to start. They had trouble with fixing the batteries to the ATV’s. They decided to go with Persy to the LOAC and MegaAres.

It was time to clean the hab for us inside and to work on the greenhab.

The afternoon was dedicated to science work. The language experiment has changed for this week and we use Keep talking and nobody explodes. This cooperative game was new and fun.

Look Ahead Plan: Start well the second week

Anomalies in work: None.

Weather: cloudy,-2°C at 7:00am, 6°C at 11:00am

Crew Physical Status: All crew members safe and sound.

EVA: LOAC and MegaAres batteries changed

Reports to be filed: Journalist report, Operations report, EVA request, Green Hab report, EVA report, Astronomy report, HSO report, Art report

Support Requested: None